# Thermodynamic & algorithmic efficiency in heterogeneous networks

### v1

@@ 3.2 Thermodynamic and algorithmic efficiency at the emergence of new C-peaks
Similar to what would be the transition from vapor/liquid to an amorphous solid, at a critical value of $r$ topological complexity increases from the trivial traits of a star-graph. Thus, structure emerges reflected by the arise of a second peak in $C$ and a roughly $C=c_{0}>0$ plateau in between the first and the last peak. Such transition can be identified by the bifurcation of the blue and green curves for $\tau^{(1)}_{peak}$ and for $\tau^{(2)}_{peak}$ respectively.
@@
Before the bifurcation the two curves overlap, and the single peak in the specific heat identifies a single resolution scale below which the informationally connected components are the single nodes, and after which the informationally connected component is the full graph. The emergence of a plateau $C=c_{0}>0$ for $r\rightarrow 2$ reflects the emergence of a progressively larger scale invariant structure. For $r\rightarrow 2$ we expect the plateau height to converge to the theoretical value $C\rightarrow c_{0}=1/2$ of the 1d-lattice [3]. The observation is cross-validated by the color code in panel (d) showing the specific heat as a function of both parameters $r$ and $\tau$. The two curves $\tau^{(1)}_{peak}$ and $\tau^{(2)}_{peak}$ are projected onto the specific heat surface in panel (d).
In panels (b), (c) the overlap between the thermodynamic efficiency, $\eta$, and and the algorithmic efficiency, $\theta$, is illustrated. The dashed lines indicate where best trade-off is obtained, while the shaded purple region indicate a tolerance region for $\theta$ defined by the intersection between the points on the efficiency curve and the standard deviation computed at the best trade-off. Striking overlap between the two efficiency is observed for the NIST and the FashionMNIST dataset, suggesting that, whether thermodynamic or algorithmic, the two objects are the same thing just viewed from different angles. Remarkably, the efficiency $\eta$ is computed at the $\tau$-resolution scales in the proximity of the transition depicted in panels (a), namely at the $\tau$ values where the bifurcation between $\tau^{(1)}_{peak}$ and $\tau^{(2)}_{peak}$ appears in panel (a).
In the seminal work [2], it was indicated to compute the thermodynamics efficiency, $\eta$, using as a reference the $\tau$-scale equal to the inverse of the smallest non-zero Laplacian eigenvalue $\lambda_{2}=1/\tau_{diff}$, as this might give the time-scale where the diffusion process has reached equilibrium and then logarithmically divide $\tau_{diff}$ to progressively define the large propagation scales, the middle scales and the small scales. Nevertheless, $\tau_{diff}$ only approximates the end of the diffusion process which is better reflected by $\tau^{(2)}_{peak}$. Panels (a) for all three figures 4, 6 and 5 show the numerical average $<\tau_{diff}>$ in red next to the $\tau^{(1,2)}_{peaks}$ curves discussed above. While the mentioned strategy is effective in the case of a transition from sparse to connected networks [2], it is less so in a transition that exclusively involves the heterogeneity of the degree distribution. We remind that the Portegy’s algorithm allows us to stay in the regime of sparsity, symbolically $\delta((N-1)/|E|-1)$.
@@
Thus, it is crucial to identify which is the $\tau$-scale involved in the phase transition from a single trivial scale, i.e. the star-graph, to more complex topological traits. Such resolution scale is the relevant scale for the computation of the efficiency $\eta$ and defines an ”ultra-violet” cut-off in the characteristic frequency of the network across the transformation, similarly to what was discussed in the context of the Laplacian-Renormalization group [3]. In section S3 of the Supplementary the efficiency, $\eta$, is computed in the range interval indicated by $<\tau_{diff}>$, i.e. the red curves depicted in panels (a). Whether computed at those scales, $\eta$ shows a change in concavity and no consistent relation with the algorithmic accuracy.
@@
Figure 4: Thermodynamic & algorithmic efficiency - NIST. In (a) the bifurcation of $\tau^{(2)}_{peak}$ and $\tau^{(2)}_{peak}$ curves identify the $\tau$-scale relevant for the transition from the single-peak to $C=c_{0}>0$ plateau reflecting the emergence of a roughly scale-invariant structure. The numerical average of the diffusion time $\tau_{diff}=<1/\lambda_{2}>$ is plotted in red for reference, see the main text for discussion. In (d) the specific heat, $C$, surface is illustrated and the $\tau^{(1,2)}_{peak}$ are projected on it. In (b)-(c) the thermodynamic efficiency, $\eta$, and the algorithmic efficiency, $\theta$, are functions of $r$. Vertical dashed-lines indicate the global maxima of the efficiencies. Remarkably, the thermodynamic efficiency is computed in the $\tau$-range of one order of magnitude from the bifurcation between $\tau^{(1)}_{peak}$ and $\tau^{(2)}_{peak}$ in (a). Note that both maxima in the efficiencies are located where a discontinuity in $\tau^{(2)}_{peak}$ is observed at $r=0.7$, far away from the transition. This is in line with the argument that indicates the NIST as a fairly simple dataset that does not benefit from the critical point. Curves are averaged over 50 independent simulations.
Figure 5: Thermodynamic & algorithmic efficiency - FashionMNIST. As for the case of the NIST, in (a) the bifurcation of $\tau^{(2)}_{peak}$ and $\tau^{(2)}_{peak}$ curves identify the $\tau$-scale relevant for the topological transition. In red the numerical average of the diffusion time $<\lambda_{2}=1/\tau_{diff}>$ is plotted for reference, see the main text for discussion. In (d) the specific heat, $C$, surface is illustrated. In (b)-(c) the efficiencies $\eta$ and $\theta$ are plotted as a function of $r$ with the dashed-lines indicating the global maxima. The $\eta$ is computed in the $\tau$-range of one order of magnitude from the $\tau$-scale indicated by the bifurcation of $\tau^{(1,2)}_{peak}$ in (a). Note that the maxima of the efficiencies is posited at the edge of the critical value of $r\sim 0.35$ indicating the FashionMNIST as a dataset more complex than the other two investigated, that does benefit from the critical region. Curves are averaged over 50 independent simulations.
Figure 6: Thermodynamic & algorithmic efficiency - MNIST. See the main text and the caption of figure 4 for a detailed explanation. The MNIST does not fully benefit from the transition point $r=0.3$, nevertheless we still observe a local maxima in the efficiency. The contraction of the $\tau^{(2)}_{peak}$ is reflected in the local maxima of $\eta$ in (b), but not in the efficiency (c). Overlap between the two efficiencies is less striking when compared to figures 4 and 5 but still valid if we include points in one standard deviation, see shaded purple area in (c), from the maximum of $\theta$. Note that the shaded area in (c) extends and includes local to global maxima in (c). Curves are averaged over 50 independent simulations.
3.3 Criticality or not to criticality
We now focus the discussion of the results obtained for the NIST and MNIST datasets. As it is shown in both panels (a) and (b) of figures 4 and 6, the ”plateau” region of the specific heat has a contraction and an expansion for $r=0.7$ for the NIST and $r=0.6$ for the MNIST. For the latter the contraction is such that the two $\tau^{(1,2)}_{peak}$ curves overlap. We mention that this effect is mitigated for larger datasets, i.e. larger networks, as we show in section S2 of the Supplementary. Moreover, for the MNIST case, the contraction of the plateau region is sufficiently strong to reflect on a local maximum in the efficiency, $\eta$, but not registered by the algorithmic efficiency $\theta$.
Note that the mentioned values of $r$ where the (largest) jump in the $\tau^{(2)}_{peak}$ is observed correspond to the $r$ values where both the best algorithmic efficiencies, $\eta$ and $\theta$, are found, as indicated by the red dashed lines in panels (c) and the multi-colored dashed lines in panels (b) (valid also for the MNIST if we include points in the one standard deviation illustrated by the shaded-purple). Nevertheless, we stress that we did not compute the efficiency at the corresponding $\tau$ scale where such discontinuity appears.
@@ 3.3 Criticality or not to criticality
Moreover, while the value of $r$ where a structural transition appears is dependent on the dataset, the $\tau\sim 10^{1}$ scale identified is the same for all the three datasets. Interestingly that $\tau$-scale is also unaffected by the number of nodes in the network (see Supplementary S3). We speculate about the possibility that this is due to the constraint on sparsity.
@@ 4 Discussion
When confronted to real world environment, in order to survive, living systems not only need the capability of interpreting the environment, but they also need the capability to quickly react with a coherent response to stimuli possibly drawn from the same distribution they adapted to [30]. Moreover the same system needs to cope with cost of sending information across its structure which can cause an enhanced exploitation of resources. In the context of machine-learning, cost can be quantified by the number of computations to be done before the satisfying answer to a task can be reported, e.g. class-matching. By using the recently developed theoretical tools of statistical physics for networks [5, 4, 31, 20, 3, 2], we have quantified diversity in the traits of the network with the von Neumann entropy and transmission flow of information as the Helmholtz free energy. As fastness of information flow and sufficiently rich representation of the outer world are two competing capabilities [2], the formation process of a networked structure is depicted as the trade-off between the two, and a variational principle in terms of thermodynamic efficiency can be devised to explain why a certain structure emerge in place of an other. We have shown that such a thermodynamic efficiency strikingly overlaps with the algorithmic efficiency defined in terms of accuracy and computational costs, suggesting that the two can be regarded as the same thing but viewed from different angles.
Notably, we have shown that the thermodynamic efficiency needs to be computed at a specific resolution scale $\tau$. We have used the specific heat $C=-dS/d\log\tau$ to identify the $\tau$-resolution involved in the transition from the trivial-structure with a single specific heat peak to emergence of more complex structures, where the specific heat shows some plateau reflecting the growth of a scale-invariant region where information is integrated at a constant rate. Such $\tau$-resolution scale is akin to an ”ultra-violet” cut-off delimiting the frequency of the network structure across the hetrogeneity-transition from star-graph to 1d-lattice. While the transition point reflects the network resolution scale where the efficiency needs to be measured, the corresponding transition value of the control parameter $r$ does not coincide with the global maximum in efficiency. We explain this with the statement that a system benefits from the critical point only when it confronts with environments of higher complexity, while it does not benefit from it when the environmental complexity is lower. This is elucidated by the use of three different datasets of increasing complexity and supported by other works in the literature [26, 10].
@@
DC: Theoretical framework, Conceptualization, Methodology, Simulations, Data curation, Software, Writing – original draft, Writing – review & editing,. LS: Writing – review & editing, Conceptualization, Supervision, Funding acquisition.
@@ Acknowledgments
The original contributions presented in the study are included in the article/supplementary material, further inquiries can be directed to the corresponding author/s. Datasets and scripts can be found at https://github.com/CipolliniDavide/LaplacianTrees.

### v2

@@ 4.1 Thermodynamic and pattern-matching efficiency at the emergence of new C-peaks
Similar to what would be the transition from vapor/liquid to an amorphous solid, at a critical value of $r$ topological complexity increases from the trivial traits of a star graph. Thus, the emergence of structure is reflected by the arise of a second peak in $C$ together with a roughly $C=c_{0}>0$ plateau in between the first and the last peak as is illustrated in panels (e)-(i). Such transition is also depicted in Figure 5 by the bifurcation of the blue and green curves for $\tau^{(1)}_{peak}$ and for $\tau^{(2)}_{peak}$ indicating the lower and upper branches in panels (g)-(i) of Figure 4.
@@
Before the bifurcation the two curves overlap and the single peak in the specific heat identifies a single trivial resolution scale below which the informationally connected components are the single nodes, and after which the informationally connected component is the full graph. The emergence of a plateau $C=c_{0}>0$ for $r\rightarrow 2$ reflects the emergence of a progressively larger scale invariant structure. For $r\rightarrow 2$ we expect the plateau height to converge to the theoretical value $C\rightarrow c_{0}=1/2$ of the 1d-lattice passing by $c_{0}=4/6$ for random trees notoriously related to the spectral dimension $d=2c_{0}$ [3]. See also Section S4 illustrating the spectral entropy and specific heat curves for the three datasets independently depicted for each $r$-value.
Remarkably, the thermodynamic efficiency is computed at the specific resolution scale $\tau\sim 10$ involved in the structural transition. Such critical resolution scale is identified by the bifurcation of the specific heat peaks depicted in panels (e)-(i) of Figure 4 and in panels (a)-(c) of Figure 5. In the seminal work [2], it is indicated to compute the thermodynamics efficiency, $\eta$, using as a reference the $\tau$-scale equal to the inverse of the Fiedler eigenvalue, i.e. the smallest non-zero Laplacian eigenvalue $\lambda_{2}=1/\tau_{diff}$, as this provides the time-scale where the diffusion process has reached equilibrium according to the authors, and then logarithmically divide $\tau_{diff}$ to progressively define the large propagation scales, the middle scales and the small scales.
@@
Nevertheless, $\tau_{diff}$ only approximates the end of the diffusion process which is better reflected by $\tau^{(2)}_{peak}$. Panels (a)-(c) of Figure 5 show the $\tau_{diff}$ in red next to the $\tau^{(1,2)}_{peaks}$ curves discussed above. The key difference is in the use of the ensemble average which employs all the eigenvalues to compute the specific heat. Thus, it is crucial to identify which is the $\tau$-scale which resolves the phase transition from the single trivial scale of the star graph to more complex topological traits of the amorphous network depicted in Figure 1. Such resolution scale is the relevant scale for the computation of the efficiency $\eta$ and defines an ”ultra-violet” cut-off in the characteristic frequency of the network across the transformation, similarly to what was discussed in the context of the Laplacian-Renormalization group [3]. In Section S3 the efficiency, $\eta$, is computed in the range interval indicated by the diffusion time as defined in [2] $\tau_{diff}=1/\lambda_{2}$, i.e. the red curves depicted in panels (a). Whether computed at those scales, $\eta$ shows a change in concavity and no consistent relation with the algorithmic accuracy.
Moreover note in Figure 5 the slow down of diffusion indicated by the $\tau^{(2)}_{peak}$ curve as a function of $r$, we remark that $\tau^{(2)}_{peak}$ is a function of the parameter $r$ only which shapes the degree distribution as the number of nodes and edges is fixed across the transition. We remind that the Portegy’s algorithm allows us to stay in the regime of sparsity, symbolically $\delta(\frac{N-1}{|E|}-1)$.
@@
Figure 4: Thermodynamic & pattern-matching efficiency. In (a)-(c) the pattern-matching efficiency, $\theta$, is plotted as a function of $r$ for the three investigated datasets of increasing complexity (see arrow at the bottom of the figure). The shaded purple area indicates the tolerance $r$-range for the optimal pattern-matching efficiency, which includes points in one standard deviation from the global maximum indicated by the vertical red-dashed line. In (d)-(f) the optimal thermodynamic efficiency, $\eta$, as a function of $r$ overlaps with the pattern-matching efficiency. Vertical dashed-lines indicate the global maxima of the efficiencies. Remarkably, the thermodynamic efficiency is measured in the $\tau$-range of one order of magnitude from the bifurcation of the specific heat, $C$, depicted in panels (g)-(i). The $\tau$-resolution scale $\tau\sim 10$ identifies the relevant length scale resolving the transition from the single-peak to the emergence of a roughly plateau region, $C=c_{0}>0$, reflecting a roughly scale-invariant structure. Note that for datasets of higher complexity the highest efficiency, both thermodynamic and algorithmic, are posited at the critical value of the structural phase transition illustrated by the specific heat in panels (g)-(i). This is in line with the argument that states that a fairly simple dataset does not benefit from the critical point. Quantities are averaged over 50 independent simulations.
Figure 5: Role of the ensemble average on the estimation of the $\tau$-scale. Panels (a)-(c) depict the $\tau^{(1,2)}_{peak}$ and $\tau_{diff}$ as a function of $r$. $\tau^{(1,2)}_{peak}$ are the $\tau$-location of the first and last peak of the specific heat in Figure 4. $\tau_{diff}$ is the Fiedler eigenvalue which provides a good estimation of the end of the diffusion process only for sufficiently large $r$. Note that the red curves retrace the profile of the upper branch of the specific heat for $\tau\gtrsim 10$ in panels (g)-(i) of Figure 4. At smaller $r$ values the equilibrium time is better reflected by $\tau^{(2)}_{peak}$. The equilibrium of the diffusion process is an increasing function of $r$. $\tau_{diff}$ curves are averaged over 50 independent simulations. The $\tau^{(1,2)}_{peak}$ curves are computed from the specific heat curves which are averaged over 50 independent simulations.
4.2 Performance at criticality
As previously discussed, the $\tau$-scale where the bifurcation in panels (e)-(i) of Figure 4 and in panels (a)-(c) of Figure 5 shows up indicates the scale involved in the structural transition and it is the resolution scale where the thermodynamic efficiency overlaps with the pattern-matching efficiency as a function of $r$. Nevertheless, the critical value of $r$ where such structural transition appears is not necessarily the value where the efficiencies have the global maximum.
To discuss this empirical result we turn our attention to the panels (a)-(c) in Figure 4 and panels (a)-(c) in Figure 5. While for the FashionMNIST dataset there is a rough correspondence between the $r\sim 0.30$ where the structural transition emerges and the $r$-range $r\in(0.30,0.50)$ where the global maximum of the efficiencies is observed, for the MNIST and NIST datasets the thermodynamic and algorithmic efficiencies show a maximum in ranges of $r$ which are farther away from the respective bifurcation points $r\sim 0.35$ and $r\sim 0.45$. The term ”far away” is quantified by the purple-shaded areas Figure 4. We rather observe in panels (b), (c), (e) and (f) of Figure 4 that the maximum efficiencies are posited where the $\tau^{(2)}_{peak}$ have a discontinuous jump in panels (b), (c) of Figure 5 at $r\sim 0.60$ for the MNIST and at $r\sim 0.70$ for the NIST. The argument holds also for the MNIST if we include $\theta$ points in the one standard deviation illustrated by the shaded-purple area. The jump for the MNIST curve in Figure 5 is such that the two $\tau^{(1,2)}_{peak}$ curves overlap. We mention that this effect is mitigated for larger datasets, i.e. larger networks, as we show in Section S2 of the Supplementary. The jump is caused by the presence of a third large peak in between $\tau^{(1)}$ and $\tau^{(2)}$ (see Figure 4 and the entropy curves depicted in Figure S6). Moreover, for the MNIST case, the jump is sufficiently strong to be reflected on a local maximum in the efficiency, $\eta$. Even though this local maximum in $\eta$ is not registered by the averaged pattern-matching efficiency $\theta$ in Figure 4(b), it is reflected in its standard deviation which allows us to include the $r$-range indicated by the shaded purple area, which corresponds to the $\eta$-peak-to-peak $r$-range in Figure 4(e).
Nevertheless, we stress once again that we did not compute the thermodynamic efficiency at the corresponding $\tau$ scale where discontinuity of the $\tau^{(2)}_{peak}$ curve appears, but we rather measured $\eta$ at the $\tau$ scale indicated by the bifurcation.
To explain the diversity in the positioning of the maximum of the efficiencies, either in the vicinity of the transition point or father away, we turn to the argument, already present in the literature [28, 10, 29], that indicates the necessity of tuning a system, or of a system to self-tune, at the border of a phase transition only weather the task requires a high level of complexity. While for the FashionMNIST, a notoriously more complex dataset when compared to both the MNIST and NIST, the optimal value for the efficiency is posited in the proximity of the critical value of $r\sim 0.30$, the optimal values of $r$ for the NIST and MNIST datasets are evidently far away from their respective transition values. The MNIST, which is the intermediate dataset in terms of complexity, is an interesting case as it shows two maxima for the thermodynamic efficiency $\eta$. While the global maximum in Figure 4(e) is far away from the critical value of $r$ illustrated by the bifurcation in panel (b) of Figure 5, the local maxima is at $r\sim 0.35-0.40$ which corresponds to the critical value of $r$ indicated by the bifurcation. This observation is in line with the hypothesis that the system takes advantage from the critical behaviour only when a fairly complex task is presented to it and in line with other observations discussed in the literature [28, 10]. Thus, it stresses the importance of dataset under analyses.
@@ 5 Discussion
When confronted to real world environment, in order to survive, living systems not only need the capability of interpreting the environment, but they also need the capability to quickly react with a coherent response to stimuli possibly drawn from the same distribution they adapted to [33]. Moreover the same system needs to cope with cost of sending information across its structure which can cause an enhanced exploitation of resources. In the context of machine-learning, cost can be quantified by the number of computations to be done before the satisfying answer to the task can be reported, e.g. class-matching. By using the recently developed theoretical tools of statistical physics of fields for networks [5, 4, 34, 21, 3, 2], we have quantified the diversity in the traits of the network with the von Neumann entropy and the transmission flow of information with the Helmholtz free energy. As fastness of information flow and sufficiently rich representation of the outer world are two competing capabilities, the formation process of a complex networked structure is interpreted as the macroscopic thermodynamic outcome of the delicate balance between the two, and a variational principle in terms of thermodynamic efficiency can be devised to explain why a certain structure emerge in place of an other [2]. We have shown that such a thermodynamic efficiency strikingly overlaps with the pattern-matching efficiency defined in terms of accuracy and computational costs, suggesting that the two can be regarded as the same thing but viewed from different angles.
Notably, we have shown that the thermodynamic efficiency needs to be computed at a specific resolution scale $\tau$. We have used the specific heat $C=-dS/d\log\tau$ to identify the $\tau$-resolution which resolves the transition from the trivial-structure, that is a single specific heat peak, to emergence of more complex structures, where the specific heat shows some plateau reflecting the growth of a scale-invariant region where information is integrated at a constant rate. Such $\tau$-resolution scale is akin to an ”ultra-violet” cut-off delimiting the frequency of the network structure across the heterogeneity-transition from star graph to 1d-lattice trough amorphous. While the transition point reflects the network resolution scale where the thermodynamic efficiency needs to be measured, the corresponding transition value of the control parameter $r$ does not coincide with the global maximum in efficiency. We explain this with the statement that a system benefits from the critical point only when it confronts with environments of higher complexity, while it does not benefit from it when the environmental complexity is lower. This is elucidated by the use of three different datasets of increasing complexity and supported by other works in the literature [28, 10, 29].
@@
DC: Theoretical framework, Conceptualization, Methodology, Simulations, Data curation, Software, Writing – original draft, Writing – review & editing. LS: Writing – review & editing, Conceptualization, Supervision, Funding acquisition.
@@ Acknowledgments
The original contributions presented in the study are included in the article/supplementary material, further inquiries can be directed to the corresponding author/s. Datasets and scripts can be found at https://github.com/CipolliniDavide/Laplacian_Trees.